# https://www.intelligence.tuc.gr/gogs/

### v2

@@ 2 Graph Neural Networks
GNNs is a class of machine learning models designed to operate on graph-structured data. Unlike traditional neural networks that process grid-structured data like images or sequences, GNNs are able to effectively handle complex relationships and dependencies inherent in graph structured data [GNN_review]. Formally, a graph is denoted as $G=(V,E)$, where $V$ represents the set of nodes (vertices) and $E$ the set of edges (connections) between nodes. In the context of GNN, a graph consists of entities which are represented as nodes and are connected by edges which correspond to relationships or interactions among them. Such structures can be found in various domains, for example in social networks, recommendation systems [recom_sys], biological networks [4700287], and more. GNNs leverage this structure to be able to analyze graph elements[powerful_xu_2019].
When operating upon graph data, different prediction tasks can be performed, which are categorized into three general types: graph-level, node-level and edge-level [GNN_review]. In a graph-level task, the objective is to predict a property or attribute for the entire graph as a whole. For example, in graph classification, the task might involve predicting the category or label of an entire graph based on the its structure and features [graphlvl, 4700287]. In node-level tasks, we want to predict a characteristic or property for each individual node within the graph. Such task is node classification, where each node is given a label or category depending on its features in the graph [nodelvl, 4700287]. Lastly, edge-level tasks focus on predicting properties or the existence of edges within the graph. This can include tasks such as link prediction, where the likelihood of a connection between pairs of nodes is predicted based on their attributes and the graph structure. Also, it can be applied on edge attribute prediction, where properties of existing edges are predicted, such as edge weights or types [edgelvl].
GNNs are able to capture the process of message exchange between nodes. At each step, nodes exchange information (messages) with their neighbors. This process allows nodes to aggregate information from their local neighborhood, enabling them to make informed decisions about their own properties or the properties of neighboring nodes [GNN_survey]. One of the most prominent features of GNNs, is their ability to incorporate contextual information from neighboring nodes [GNN_review]. For example, in a node classification task, each node can aggregate information from its neighboring nodes to refine its own—and possibly not so far known—classification label.
@@ 2.1 GNN variations
While GNNs offer powerful capabilities for learning from graph-structured data, researchers have developed a number of variants to address specific challenges that exist in the different application domains. Some notable variants of GNNs include GCNs [GCN], GATs [gat] and GraphSAGE [graphsage].
GCNs generalize the concept of convolutional neural networks to graphs, enabling effective information propagation and feature learning across graph nodes. They have been successfully applied in tasks such as node classification and graph classification [GCN]. At the heart of GCNs lies the notion of graph convolution, which enables nodes to aggregate and propagate information from their neighbors. Mathematically, the graph convolution operation in GCN can be expressed as follows:
@@
where $H^{(l)}$ represents the node feature matrix at layer $l$ for every single node, $\widetilde{A}$ is the adjacency matrix of the graph with self-loops added, $\widetilde{D}$ is the corresponding degree matrix, which is a diagonal matrix with the summation row-wise of the amount of connections per node, $W^{(l)}$ is the learnable weight matrix for layer $l$ used to linearly transform the node features, and $\sigma$ denotes the activation function such as ReLU [GCN], applied element-wise to the result of the linear transformation.
Figure 1: GCN architecture [GCN]
@@
To illustrate the application of GCNs, consider the task of node classification in a social network. Each node in the graph represents a user, and the edges denote relationships between users (e.g., friendship). By leveraging GCNs, we can learn representations of users based on their interactions and predict user attributes or labels. For example, we can predict the demographic characteristics of users (e.g., age, gender) based on their network connections and activity patterns. Furthermore, GCNs are successfully applied for graph classification tasks, where the goal is to classify entire graphs based on their structural properties. For instance, in molecular graph analysis each graph represents a chemical compound, and the nodes and edges correspond to atoms and bonds, respectively. GCNs can effectively capture the molecular structure and predict properties of compounds, such as toxicity or drug activity [GNN_review]. In summary, GCNs provide a powerful framework for learning from graph-structured data, enabling tasks such as node classification and graph classification. By extending convolutional operations to graphs, GCNs facilitate information propagation and feature learning across nodes, making them suitable for a wide range of applications in domains such as social networks, biology, and recommendation systems.
Another noteworthy extension of GNNs are GATs [gat, GNN_review], which emerge as a neural network architecture for processing graph-structured data. Whether it is for modeling social networks, molecular structures, or language syntax, GATs leverage attention mechanisms to weight the importance of the different nodes and edges in a graph. GATs introduce attention mechanisms into the message passing process, allowing nodes to dynamically weigh the importance of information received from their neighbors [attention_vaswani_2017]. By assigning attention coefficients to edges, GATs can capture complex relationships and dependencies in graph data more effectively [gat].
In a GAT, each edge connecting two nodes is associated with an attention score. These scores, denoted as $\alpha_{ij}$ for the edge between node $i$ and node $j$, are computed based on a learned weight matrix $W$, which is trained alongside the rest of the neural network. The attention scores $\alpha_{ij}$ reflect the importance or relevance of node $j$ to node $i$ when passing messages [MA_GAT, gat]. The computation of attention score is typically performed using a form of attention mechanism, such as the softmax function, which ensures that the attention scores sum to one across all neighboring nodes:
@@
Here, $h_{i}$ and $h_{j}$ are the feature representations of nodes $i$ and $j$, respectively, and $a$ is a learnable attention parameter vector [gat], the combination of the mentioned elements is also presented in Figure 2.
Figure 2: Computation of attention coefficients [gat]
@@
This attention mechanism allows to concentrate on the most relevant information, thus facilitating more informed and accurate predictions. The versatility of GATs is underscored by their application in diverse tasks such as node classification, link prediction, and graph generation [GNN_review]. GATs address the limitations of traditional GCNs by enabling nodes to adaptively focus on influential neighbors. This attention mechanism enhances context-aware feature propagation and precise predictions within graphs. They can also be applied in multi-agent settings, due to the inherent ability of GATs to capture intricate dependencies within graph structures. This allows them to capture evolving relationships between the nodes [MA_GAT].
@@
To avoid confusion, it is important to highlight some of the differences between GCNs and GATs. GCNs represent a class of neural networks tailored for graph structures, employing convolutional operations to relay information among neighboring nodes within a graph. This mechanism, allows GCNs to extract informative features representing the graph’s topology, which is useful in tasks such as node classification and link prediction [GCN, GNN_review]. In contrast, GATs use attention mechanisms to dynamically assign importance to adjacent nodes during information propagation. This flexibility enables GATs to learn varying degrees of significance among neighbors, offering a more nuanced approach compared to the fixed weighting employed by GCNs. Notably, GATs have exhibited superior performance over GCNs across various benchmark datasets, particularly in scenarios characterized by sparse graphs or tasks requiring the learning of extensive dependencies [gat, GNN_survey].
@@
GraphSAGE (Graph Sample and Aggregation) is a framework for inductive representation learning on large graphs. It samples and aggregates information from node neighborhoods, enabling scalable and efficient learning on graphs with millions or even billions of nodes [graphsage]. Existing embedding methods focus on fixed graphs and use matrix-factorization for each node, lacking the ability to generalize to unseen data. In contrast, GraphSAGE uses the node features in order to develop an embedding function applicable to nodes not present in training. Instead of treating each node in isolation, GraphSAGE considers the collective information from a node’s neighbors. For example, in a social network, we are interested in learning more about a certain node, in order to get the full picture we also check information about its connections and their characteristics. In a similar manner, to understand the context of a node, GraphSAGE aggregates information from the nodes in the neighborhood of our interest. To achieve this, GraphSAGE trains a set of aggregator functions. The intuition is that at each training iteration these functions learn to combine information from neighboring nodes at different depths, as shown in Figure 3.
Figure 3: GraphSAGE architecture [graphsage]
@@
The aggregated representation of node $v$ at layer $k$ is represented by the term $h^{(k)}_{v}$ and captures the information from the neighboring nodes up to layer $k$. Respectively, $h^{(k-1)}_{v}$ represents the feature vector of a neighboring node $u$ at the previous layer k1k-1. The set of neighboring nodes of node $v$ is denoted by $N(v)$. The component $AGGREGATE_{k}$ is the aggregation function used to aggregate information from the neighboring nodes. It operates on the feature vectors of neighboring nodes $h^{(k-1)}_{v}$ and produces the aggregated representation $h^{(k)}_{v}$ for node $v$. For each $v$, we gather the feature vectors $h^{(k-1)}_{v}$ of all its neighboring nodes $u$ at the previous layer k1k-1. These feature vectors are then aggregated using $AGGREGATE_{k}$, which computes a summary representation of the neighboring nodes’ features, capturing the local neighborhood information of node $v$. The result $h^{(k)}_{v}$ captures the information for the neighborhood of node $v$ up to layer $k$, which is then used as input for the next layer. One of GraphSAGE’s prominent features is its ability to be employed in different types of graphs, apart from social networks. It can also be used, for example, to understand the connections between different proteins in a biological network [PPI]. GraphSAGE’s performance is mainly tested in node-classification tasks, such as citation networks, where it is able e.g., to predict the topic of a paper by observing only its citations and its abstract [graphsage].
@@
There are also GNN variants that operate on more complex graphs, such as on multigraphs [harary1969graph] or hypergraphs [berge1973graphs, christianos2017efficient]. For instance, hypergraphs can come in handy for representing settings where communities of nodes share similar characteristics and can be thus seen as belonging in a hyperedge. How to design and train GNNs on such complex graphs is an active area of research [DBLP:conf/nips/YadatiNYNLT19, zhong2023hierarchical].
Another important aspect is the underlying uncertainty in some cases of graph structures. Most of the examined methods are not able to capture uncertainty in case this is inherent in some problem settings. An approach that is able to handle such issues though is Bayesian Graph Convolutional Neural Networks [zhang2019bayesian]. In Bayesian Neural Networks, the internal network weights are considered as random variables, a fact that renders the NN’s output to be a random variable as well. Accordingly, the prediction for a new input $x$ is calculated by integrating with respect to the posterior distribution of the weights $W$:
@@
where $\mathbf{Y}_{\mathcal{L}}$ are the observed node labels, $\mathcal{G}_{obs}$ the observed graph structure, and $\lambda$ parameters that characterize a family of random graphs. Since this integral is intractable, approximation methods can be used for its calculation, such as variational inference and Monte Carlo Markov Chains. Bayesian GNNs have been used in a plethora of domains, such as multiple instance learning [pal2022bag], network analysis, biochemistry, recommendation systems [9555949], as well as in non-parametric graph distribution cases [pmlr-v124-pal20a].
@@ 2.2 GNN limitations
GNNs can become intricate and challenging to interpret, especially when multiple layers and attention mechanisms are incorporated [attention_vaswani_2017, gat]. Additionally, their performance is contingent on hyperparameter settings, including the number of layers, attention mechanisms, and learning rates. GNNs also face challenges in scenarios where limited historical data is available for new or rarely encountered coalition formations. Furthermore, temporal dynamics is another element that GNNs face trouble with, as GNNs inherently assume a static graphic structure and may struggle to capture evolving relationships over time. Overall, GNNs can also be applied—though with appropriate calibration—to multi-agent systems, which inherently possess graph-like structures, where agents and their relationships can be represented as nodes and edges in a graph [GNN_review].
@@ 2.3 GNN applications in Multiagent settings
An important work highlighting the use of GNNs in MA environments is that of [MA_GAT]. The paper proposes a novel game abstraction mechanism based on two-stage attention network called G2ANet, which models the relationship between agents using a complete graph, to detect interactions and their importance. The paper addresses the problem of policy learning in large-scale multi-agent systems, where the large number of agents and complex game relationships make the learning process challenging. In order to capture interactions between agents, existing methods use pre-defined rules, which are not suitable for large-scale environments due to the complexity of transforming those interactions into rules. The integration of G2ANet into GNN-based MARL is able to simplify the learning process as well as improve asymptotic performance in comparison to state-of-the-art algorithms [MAAC, MADDPG, IC3Net].
@@
Another noteworthy employment of GNNs in a multiagent setting is that in the work of [DICG]. The Deep Implicit Coordination Graph (DICG) architecture is introduced, with the goal of enhancing coordination in multiagent reinforcement learning scenarios. The DICG consists of a dynamic coordination graph inference module accompanied by a GNN module. In such scenarios, coordination graphs are usually computed based on factorized representations, which leads to challenges in both scalability and adaptability. However, DICG employs attention mechanisms to learn the agent observation-dependent coordination graph with soft weights [DICG]. In that way, they avoid relying on domain-specific heuristics and are able to represent complex relationships between agents across various multiagent domains.
@@
In the above context, the approach used in [DICG] follows the principles of the CPA but not the SIH. This is because it focuses on improving the ability of the agents to learn coordination strategies considering collective goals, by inferring dynamic coordination graph structures. Overall, the DICG architecture offers a flexible framework for multiagent reinforcement learning, leveraging attention mechanisms and graph neural networks to infer coordination structures. It improves coordination efficiency in domains with a large number of agents. However, it has not yet been applied into off-policy learning and its sampling efficiency shows grounds for improvement in the future. Similar to [DICG] are the methods implemented in [GCRL_MA] and [GNN_MARL_DICGLIKE].
@@
The work of [gatdrl] is a combination of GNNs in MA settings in the context of resource allocation in dense cellular networks. Their approach focuses on the context of network slicing and aims to address the challenge of efficiently allocating resources to different slices, such as devices and applications, while maximizing the network performance and user satisfaction. In order to tackle this problem, they leverage GATs and MARL. By utilizing the attention mechanisms of GATs, they can effectively capture the interactions between different network entities and organize relevant information. This way, the complex relationships and dependencies among network elements are effectively modeled, including base stations, users and network slices. The employment of MARL techniques enables collaborative decision-making among multiple agents, such as base stations, in dynamically allocating resources based on network conditions and user demands. Such a collaborative approach is able to enhance the adaptability in resource management leading to improved network performance as well as resource utilization. The proposed approach has broad applications in various scenarios where efficient resource management is crucial, including next-generation cellular networks, edge computing environments, and Internet of Things (IoT) networks. However, it is limited by challenges regarding computational complexity, scalability, and generalization to diverse network environments. Overall, [gatdrl] presents an innovative solution to address resource management challenges in dense cellular networks using GATs and MARL. Further research and experimentation is required to validate its effectiveness and address potential limitations in real-world deployments. Other works in the same context are those of [marl-gq] and [magnet]. Interestingly, this paper does not rely on the CPA, since the coordination mechanism implemented with the use of GATs, is to capture the fluctuating service demands on the network. This does not include any priors that must be updated, but instead tries in real-time to achieve dynamic collaboration among the base stations. Additionally, we could not claim that it is based on the SIH, due to the fact that the goal of the approach is the optimization of the resource allocation strategy, which is a collective goal rather than having each base station try to maximize some individual reward.
@@
Lastly, we refer to the work of [marl-autdriving]. This study explores the domain of autonomous vehicle control, utilizing RL alongside latent state estimation and spatiotemporal correlations. The proposed framework integrates latent estimation with relational representation, introducing an innovative strategy to improve policy efficiency in autonomous driving situations. The paper introduces a reinforcement learning framework that explicitly seeks to infer the hidden states of adjacent drivers by utilizing GNNs, thereby notably enhancing the performance of autonomous driving policies. It offers designations for hidden states and impact transmission structures, consequently improving the interpretability within autonomous driving situations. The approach integrates hidden state inference and relational modeling, with the objective of enhancing both the success rates of navigation and the accuracy of hidden state inference within intricate driving situations. By employing GNNs as in the proposed STGSage model, we can attain a deep comprehension of spatial relationships and temporal advancements among different traffic entities, thereby enhancing decision-making in intricate driving scenarios. The results exhibit a significant improvement in the rates of successful navigation, especially in intricate T-intersection situations. Through the integration of latent inference and RL, the model effectively tackles the obstacles related to interpretability and efficacy, presenting a hopeful resolution for autonomous driving mechanisms. The unique strategy not only enhances the accuracy of navigation results but also creates a solid basis for future advancements in the field of autonomous driving technology. Despite its advancements, the study acknowledges certain limitations, notably the complexity introduced by varying numbers of surrounding vehicles in different scenarios. However, the proposed framework showcases a substantial improvement in latent inference accuracy, particularly in T-intersection scenarios, underscoring its potential for real-world application in autonomous driving contexts.
@@
The work of [marl-autdriving] focuses on using RL techniques to enable autonomous vehicles to navigate complex driving scenarios independently, without relying on a shared understanding or coordination among agents. Thus, the approach does not rely on the CPA. However, it does follow the SIH, since each vehicle learns its own policy without direct coordination with other vehicles. Also, each vehicle is focused on optimizing its own behavior to achieve its objectives, such as reaching its destination safely and efficiently.
@@
GCN [GCN] GAT [gat] GraphSAGE [graphsage] G2ANet [MA_GAT] DICG [DICG] GAT-based MARL [gatdrl] STGSage[marl-autdriving]
@@
Table 1: A taxonomy of the GNN-based algorithms that we cover in our review.

### v3

@@ 2 Graph Neural Networks
GNNs is a class of machine learning models designed to operate on graph-structured data. Unlike traditional neural networks (NNs) that process grid-structured data like images or sequences, GNNs are able to effectively handle complex relationships and dependencies inherent in graph structured data [164]. Formally, a graph is denoted as $G=(V,E)$, where $V$ represents the set of nodes (vertices) and $E$ the set of edges (connections) between nodes. In the context of GNN, a graph consists of entities which are represented as nodes and are connected by edges which correspond to relationships or interactions among them. Such structures can be found in various domains, for example in social networks, recommendation systems [38], biological networks [115], and more. GNNs leverage this structure to be able to analyze graph elements [153].
When operating upon graph data, different prediction tasks can be performed, which are categorized into three general types: graph-level, node-level and edge-level [164]. In a graph-level task, the objective is to predict a property or attribute for the entire graph as a whole. For example, in graph classification, the task might involve predicting the category or label of an entire graph based on the its structure and features [37, 115]. In node-level tasks, we want to predict a characteristic or property for each individual node within the graph. Such task is node classification, where each node is given a label or category depending on its features in the graph [4, 115]. Lastly, edge-level tasks focus on predicting properties or the existence of edges within the graph. This can include tasks such as link prediction, where the likelihood of a connection between pairs of nodes is predicted based on their attributes and the graph structure. Also, it can be applied on edge attribute prediction, where properties of existing edges are predicted, such as edge weights or types [160].
GNNs are able to capture the process of message exchange between nodes. At each step, nodes exchange information (messages) with their neighbors. This process allows nodes to aggregate information from their local neighborhood, enabling them to make informed decisions about their own properties or the properties of neighboring nodes [150]. One of the most prominent features of GNNs, is their ability to incorporate contextual information from neighboring nodes [164]. For example, in a node classification task, each node can aggregate information from its neighboring nodes to refine its own—and possibly not so far known—classification label.
@@ 2.1 GNN variations
While GNNs offer powerful capabilities for learning from graph-structured data, researchers have developed a number of variants to address specific challenges that exist in the different application domains. Some notable variants of GNNs include GCNs [69], GATs [141] and GraphSAGE [53].
GCNs generalize the concept of convolutional neural networks to graphs, enabling effective information propagation and feature learning across graph nodes. They have been successfully applied in tasks such as node classification and graph classification [69]. At the heart of GCNs lies the notion of graph convolution, which enables nodes to aggregate and propagate information from their neighbors. Mathematically, the graph convolution operation in GCN can be expressed as follows:
@@
where $H^{(l)}$ represents the node feature matrix at layer $l$ for every single node, $\widetilde{A}$ is the adjacency matrix of the graph with self-loops added, $\widetilde{D}$ is the corresponding degree matrix, which is a diagonal matrix with the summation row-wise of the amount of connections per node, $W^{(l)}$ is the learnable weight matrix for layer $l$ used to linearly transform the node features, and $\sigma$ denotes the activation function such as ReLU [69], applied element-wise to the result of the linear transformation.
Figure 1: GCN architecture [69]
@@
To illustrate the application of GCNs, consider the task of node classification in a social network. Each node in the graph represents a user, and the edges denote relationships between users (e.g., friendship). By leveraging GCNs, we can learn representations of users based on their interactions and predict user attributes or labels. For example, we can predict the demographic characteristics of users (e.g., age, gender) based on their network connections and activity patterns. Furthermore, GCNs are successfully applied for graph classification tasks, where the goal is to classify entire graphs based on their structural properties. For instance, in molecular graph analysis each graph represents a chemical compound, and the nodes and edges correspond to atoms and bonds, respectively. GCNs can effectively capture the molecular structure and predict properties of compounds, such as toxicity or drug activity [164]. In summary, GCNs provide a powerful framework for learning from graph-structured data, enabling tasks such as node classification and graph classification. By extending convolutional operations to graphs, GCNs facilitate information propagation and feature learning across nodes, making them suitable for a wide range of applications in domains such as social networks, biology, and recommendation systems.
Another noteworthy extension of GNNs are GATs [141, 164], which emerge as a neural network architecture for processing graph-structured data. Whether it is for modeling social networks, molecular structures, or language syntax, GATs leverage attention mechanisms to weight the importance of the different nodes and edges in a graph. GATs introduce attention mechanisms into the message passing process, allowing nodes to dynamically weigh the importance of information received from their neighbors [140]. By assigning attention coefficients to edges, GATs can capture complex relationships and dependencies in graph data more effectively [141].
In a GAT, each edge connecting two nodes is associated with an attention score. These scores, denoted as $\alpha_{ij}$ for the edge between node $i$ and node $j$, are computed based on a learned weight matrix $W$, which is trained alongside the rest of the neural network. The attention scores $\alpha_{ij}$ reflect the importance or relevance of node $j$ to node $i$ when passing messages [80, 141]. The computation of attention score is typically performed using a form of attention mechanism, such as the softmax function, which ensures that the attention scores sum to one across all neighboring nodes:
@@
Here, $h_{i}$ and $h_{j}$ are the feature representations of nodes $i$ and $j$, respectively, and $a$ is a learnable attention parameter vector [141], the combination of the mentioned elements is also presented in Figure 2.
Figure 2: Computation of attention coefficients [141]
@@
This attention mechanism allows to concentrate on the most relevant information, thus facilitating more informed and accurate predictions. The versatility of GATs is underscored by their application in diverse tasks such as node classification, link prediction, and graph generation [164]. GATs address the limitations of traditional GCNs by enabling nodes to adaptively focus on influential neighbors. This attention mechanism enhances context-aware feature propagation and precise predictions within graphs. They can also be applied in multi-agent settings, due to the inherent ability of GATs to capture intricate dependencies within graph structures. This allows them to capture evolving relationships between the nodes [80].
@@
To avoid confusion, it is important to highlight some of the differences between GCNs and GATs. GCNs represent a class of neural networks tailored for graph structures, employing convolutional operations to relay information among neighboring nodes within a graph. This mechanism, allows GCNs to extract informative features representing the graph’s topology, which is useful in tasks such as node classification and link prediction [69, 164]. In contrast, GATs use attention mechanisms to dynamically assign importance to adjacent nodes during information propagation. This flexibility enables GATs to learn varying degrees of significance among neighbors, offering a more nuanced approach compared to the fixed weighting employed by GCNs. Notably, GATs have exhibited superior performance over GCNs across various benchmark datasets, particularly in scenarios characterized by sparse graphs or tasks requiring the learning of extensive dependencies [141, 150].
@@
GraphSAGE (Graph Sample and Aggregation) is a framework for inductive representation learning on large graphs. It samples and aggregates information from node neighborhoods, enabling scalable and efficient learning on graphs with millions or even billions of nodes [53]. Existing embedding methods focus on fixed graphs and use matrix-factorization for each node, lacking the ability to generalize to unseen data. In contrast, GraphSAGE uses the node features in order to develop an embedding function applicable to nodes not present in training. Instead of treating each node in isolation, GraphSAGE considers the collective information from a node’s neighbors. For example, in a social network, we are interested in learning more about a certain node, in order to get the full picture we also check information about its connections and their characteristics. In a similar manner, to understand the context of a node, GraphSAGE aggregates information from the nodes in the neighborhood of our interest. To achieve this, GraphSAGE trains a set of aggregator functions. The intuition is that at each training iteration these functions learn to combine information from neighboring nodes at different depths, as shown in Figure 3.
Figure 3: GraphSAGE architecture [53]
@@
The aggregated representation of node $v$ at layer $k$ is represented by the term $h^{(k)}_{v}$ and captures the information from the neighboring nodes up to layer $k$. Respectively, $h^{(k-1)}_{v}$ represents the feature vector of a neighboring node $u$ at the previous layer k1k-1. The set of neighboring nodes of node $v$ is denoted by $N(v)$. The component $AGGREGATE_{k}$ is the aggregation function used to aggregate information from the neighboring nodes. It operates on the feature vectors of neighboring nodes $h^{(k-1)}_{v}$ and produces the aggregated representation $h^{(k)}_{v}$ for node $v$. For each $v$, we gather the feature vectors $h^{(k-1)}_{v}$ of all its neighboring nodes $u$ at the previous layer k1k-1. These feature vectors are then aggregated using $AGGREGATE_{k}$, which computes a summary representation of the neighboring nodes’ features, capturing the local neighborhood information of node $v$. The result $h^{(k)}_{v}$ captures the information for the neighborhood of node $v$ up to layer $k$, which is then used as input for the next layer. One of GraphSAGE’s prominent features is its ability to be employed in different types of graphs, apart from social networks. It can also be used, for example, to understand the connections between different proteins in a biological network [166]. GraphSAGE’s performance is mainly tested in node-classification tasks, such as citation networks, where it is able e.g., to predict the topic of a paper by observing only its citations and its abstract [53].
@@
There are also GNN variants that operate on more complex graphs, such as on multigraphs [54] or hypergraphs [12, 27]. For instance, hypergraphs can come in handy for representing settings where communities of nodes share similar characteristics and can be thus seen as belonging in a hyperedge. How to design and train GNNs on such complex graphs is an active area of research [154, 163].
Another important aspect is the underlying uncertainty in some cases of graph structures. Most of the examined methods are not able to capture uncertainty in case this is inherent in some problem settings. An approach that is able to handle such issues though is Bayesian Graph Convolutional Neural Networks [161]. In Bayesian Neural Networks, the internal network weights are considered as random variables, a fact that renders the neural NN’s output to be a random variable as well. Accordingly, the prediction for a new input $x$ is calculated by integrating with respect to the posterior distribution of the weights $W$:
@@
where $\mathbf{Y}_{\mathcal{L}}$ are the observed node labels, $\mathcal{G}_{obs}$ the observed graph structure, and $\lambda$ parameters that characterize a family of random graphs. Since this integral is intractable, approximation methods can be used for its calculation, such as variational inference and Monte Carlo Markov Chains. Bayesian GNNs have been used in a plethora of domains, such as multiple instance learning [104], network analysis, biochemistry, recommendation systems [123], as well as in non-parametric graph distribution cases [103].
@@ 2.2 GNN limitations
GNNs can become intricate and challenging to interpret, especially when multiple layers and attention mechanisms are incorporated [140, 141]. Additionally, their performance is contingent on hyperparameter settings, including the number of layers, attention mechanisms, and learning rates. GNNs also face challenges in scenarios where limited historical data is available for new or rarely encountered coalition formations. Furthermore, temporal dynamics is another element that GNNs face trouble with, as GNNs inherently assume a static graphic structure and may struggle to capture evolving relationships over time. Overall, GNNs can also be applied—though with appropriate calibration—to multi-agent systems, which inherently possess graph-like structures, where agents and their relationships can be represented as nodes and edges in a graph [164].
@@ 2.3 GNN applications in Multiagent settings
An important work highlighting the use of GNNs in MA environments is that of [80]. The paper proposes a novel game abstraction mechanism based on two-stage attention network called G2ANet, which models the relationship between agents using a complete graph, to detect interactions and their importance. The paper addresses the problem of policy learning in large-scale multi-agent systems, where the large number of agents and complex game relationships make the learning process challenging. In order to capture interactions between agents, existing methods use pre-defined rules, which are not suitable for large-scale environments due to the complexity of transforming those interactions into rules. The integration of G2ANet into GNN-based MARL is able to simplify the learning process as well as improve asymptotic performance in comparison to state-of-the-art algorithms [60, 82, 128].
@@
Another noteworthy employment of GNNs in a multiagent setting is that in the work of [75]. The Deep Implicit Coordination Graph (DICG) architecture is introduced, with the goal of enhancing coordination in multiagent reinforcement learning scenarios. The DICG consists of a dynamic coordination graph inference module accompanied by a GNN module. In such scenarios, coordination graphs are usually computed based on factorized representations, which leads to challenges in both scalability and adaptability. However, DICG employs attention mechanisms to learn the agent observation-dependent coordination graph with soft weights [75]. In that way, they avoid relying on domain-specific heuristics and are able to represent complex relationships between agents across various multiagent domains.
@@
In the above context, the approach used in [75] follows the principles of the CPA but not the SIH. This is because it focuses on improving the ability of the agents to learn coordination strategies considering collective goals, by inferring dynamic coordination graph structures. Overall, the DICG architecture offers a flexible framework for multiagent reinforcement learning, leveraging attention mechanisms and graph neural networks to infer coordination structures. It improves coordination efficiency in domains with a large number of agents. However, it has not yet been applied into off-policy learning and its sampling efficiency shows grounds for improvement in the future. Similar to [75] are the methods implemented in [62] and [92].
@@
The work of [119] is a combination of GNNs in MA settings in the context of resource allocation in dense cellular networks. Their approach focuses on the context of network slicing and aims to address the challenge of efficiently allocating resources to different slices, such as devices and applications, while maximizing the network performance and user satisfaction. In order to tackle this problem, they leverage GATs and MARL. By utilizing the attention mechanisms of GATs, they can effectively capture the interactions between different network entities and organize relevant information. This way, the complex relationships and dependencies among network elements are effectively modeled, including base stations, users and network slices. The employment of MARL techniques enables collaborative decision-making among multiple agents, such as base stations, in dynamically allocating resources based on network conditions and user demands. Such a collaborative approach is able to enhance the adaptability in resource management leading to improved network performance as well as resource utilization. The proposed approach has broad applications in various scenarios where efficient resource management is crucial, including next-generation cellular networks, edge computing environments, and Internet of Things (IoT) networks. However, it is limited by challenges regarding computational complexity, scalability, and generalization to diverse network environments. Overall, [119] presents an innovative solution to address resource management challenges in dense cellular networks using GATs and MARL. Further research and experimentation is required to validate its effectiveness and address potential limitations in real-world deployments. Other works in the same context are those of [16] and [86]. Interestingly, this paper does not rely on the CPA, since the coordination mechanism implemented with the use of GATs, is to capture the fluctuating service demands on the network. This does not include any priors that must be updated, but instead tries in real-time to achieve dynamic collaboration among the base stations. Additionally, the approach is arguably not based on the SIH, due to the fact that the goal of the approach is the optimization of the resource allocation strategy, which is a collective goal rather than having each base station try to maximize some individual reward.
@@
The study presented in [85] explores the domain of autonomous vehicle control, utilizing RL alongside latent state estimation and spatiotemporal correlations. The proposed framework integrates latent estimation with relational representation, introducing an innovative strategy to improve policy efficiency in autonomous driving situations. The paper introduces a reinforcement learning framework that explicitly seeks to infer the hidden states of adjacent drivers by utilizing GNNs, thereby notably enhancing the performance of autonomous driving policies. It offers designations for hidden states and impact transmission structures, consequently improving the interpretability within autonomous driving situations. The approach integrates hidden state inference and relational modeling, with the objective of enhancing both the success rates of navigation and the accuracy of hidden state inference within intricate driving situations. By employing GNNs as in the proposed STGSage model, we can attain a deep comprehension of spatial relationships and temporal advancements among different traffic entities, thereby enhancing decision-making in intricate driving scenarios. The results exhibit a significant improvement in the rates of successful navigation, especially in intricate T-intersection situations. Through the integration of latent inference and RL, the model effectively tackles the obstacles related to interpretability and efficacy, presenting a hopeful resolution for autonomous driving mechanisms. The unique strategy not only enhances the accuracy of navigation results but also creates a solid basis for future advancements in the field of autonomous driving technology. Despite its advancements, the study acknowledges certain limitations, notably the complexity introduced by varying numbers of surrounding vehicles in different scenarios. However, the proposed framework showcases a substantial improvement in latent inference accuracy, particularly in T-intersection scenarios, underscoring its potential for real-world application in autonomous driving contexts.
@@
The work of [85] focuses on using RL techniques to enable autonomous vehicles to navigate complex driving scenarios independently, without relying on a shared understanding or coordination among agents. Thus, the approach does not rely on the CPA. However, it does follow the SIH, since each vehicle learns its own policy without direct coordination with other vehicles. Also, each vehicle is focused on optimizing its own behavior to achieve its objectives, such as reaching its destination safely and efficiently.
Koresis et al. in [71] propose harnessing the power of Hypergraph Neural Networks (HGNNs) that fit generic hypergraph-structured historical representations of coalitional task executions, in order to learn the unknown values of coalitional configurations undertaking the tasks. Unlike previous works in the literature, here the coalitional value uncertainty is due to uncertainty regarding the values of synergies among the different agent types: the unknown values of the agent types’ synergies in completing various coalitional tasks are what is learned by the HGNN. However, the fitted model by itself cannot be used to provide suggestions on which coalitions to form; it can only be queried for the values of given coalition–task configurations. To actually provide coalitional suggestions, the approach relies on informed search approaches that incorporate HGNN output as an indicator of the quality of the proposed coalition configurations.
Finally, the work of [109] presents a GNN-based approach to learn the models of agents in open ad-hoc teamwork settings. In such settings, a learning agent is tasked with efficiently collaborating with an unknown group of teammates whose composition may change over time, since agents of different types can dynamically enter and leave the team. Specifically, the learner undertakes the task to fit the behavioral models of participating agents and to estimate the optimal policy in an open ad hoc teamwork setting modeled as a stochastic Bayesian game. The proposed Graph-based Policy Learning (GPL) approach puts forward an action-value computation that integrates GNN-learned agent and joint-action value models to produce action-value estimates. In some detail, the method utilizes coordination graphs to recognize the other agents’ actions effect on the learner, and then GNN to predict the upcoming actions. Then, [108] extends the aforementioned framework to also account for the more realistic case of partial state observability—providing new algorithms specifically designed for such partially observable settings, alongside a novel formal model for open ad hoc teamwork.
@@
GCN [69] GAT [141] GraphSAGE [53] G2ANet [80] DICG [75] GAT-based MARL [119] STGSage [85] HGNN [71] GPL [108]
@@
Table 1: A taxonomy of the GNN-based algorithms that we cover in our review. The spectral domain is where graph features are represented in the graph Fourier basis, so “convolutions” are defined as filtering by functions of Laplacian eigenvalues. The spatial domain is where graph convolutions are defined directly on the graph structure by aggregating and transforming information from a node’s local neighborhood (message passing) in the vertex/node domain.